OpenStudy (minimato):

1. You have 2 m3 of a mineral ore sample. It has a mass of 1 × 104 kg. Calculate the density of the ore sample and express your results in g/cm3 with appropriate scientific notation

OpenStudy (minimato):

I'm not exactly sure if I'm going in the right direction, but here's my thought process. D = m/V (m is in grams and V is in cm^3) 2 m^3 = 200 cm^3 (is that right?) 1 x 10^4 kg = 10,000 kg (right?) 10,000 kg = 10,000,000 g (right?) 10,000,000 g/200 cm^3 = 50,000 g/cm^3 Is that the right answer? I may have made some miscalculations but I'm not sure...

OpenStudy (minimato):

Oh yeah, and 50,000 g/cm^3 would be 5 g/cm^3 x 10^4 in it's scientific notation form, right?

OpenStudy (abb0t):

Density is defined as: \(\sf \color{red}{\frac{mass}{volume}}\)

OpenStudy (minimato):

I know. Can you tell me if I solved the problem correctly or if I strayed wrong somewhere?